OpenStudy (anonymous):

Find the point on the line 6x+y=9 that is closest to the point (-3,1).

OpenStudy (anonymous):

6x + y = 9 y = -6x + 9 slope is -6 perpendicular would have slope 1/6 (y - 1) = (1/6) (x + 3) y - 1 = (1/6) x + (1/2) multiplying both sides by 6 6y - 6 = x + 3 x - 6y = -9 _______1 From above we know given equation is y = -6x + 9 substitute this equation in 1 x - 6(-6x + 9) = -9 x + 36x - 54 = -9 37 x = 45 x = 45/37 y = -6(45/37) + 9 y = -270/37 + (333/37) y = 63/37 point is (45/37, 63/37)

OpenStudy (anonymous):

@ASAAD123 http://answers.yahoo.com/question/index?qid=20081106202035AAlEt5X The above solution is algebraic, but since this is a calculus section lets explore calculus type solutions: